OpenStudy (queenechinose):

Is the derivative of c/ax+b equal to -ca/(ax+b)^2?

OpenStudy (santhosh):

Yaa you r ri8

OpenStudy (trojanpoem):

c ( a x + b )^ -1 = - c(ax+b)^-2 ( a ) = -ca/(ax+b)^2 correct